OpenStudy (anonymous):

What is the geometric mean of 5 and 15?

OpenStudy (asnaseer):

geometric mean of two numbers is the square root of their product.

OpenStudy (asnaseer):

do you know how to calculate it for the numbers you were given?

OpenStudy (anonymous):

Thank you for helping. i understand now

OpenStudy (anonymous):

I like using proportions to see this definition 5/x = x/15 or x^2 = 75